# Myers–Steenrod theorems for metric and singular Riemannian foliations

### v1

@@
Abstract.
We prove that the group of isometries preserving a metric foliation on a closed Alexandrov space $X$, or a singular Riemannian foliation on a manifold $M$ is a closed subgroup of the isometry group of $X$ in the case of a metric foliation, or of the isometry group of $M$ for the case of a singular Riemannian foliation. We obtain a sharp upper bound for the dimension of these subgroups and show that, when equality holds, the foliations that realize this upper bound are induced by fiber bundles whose fibers are round spheres or projective spaces. Moreover, singular Riemannian foliations that realize the upper bound are induced by smooth fiber bundles whose fibers are round spheres or projective spaces.
Key words and phrases:
Alexandrov space, submetry, isometry group, singular Riemannian foliation, Lie group
2020 Mathematics Subject Classification:
53C12,53C20,53C21,53C23,53C24,51K10
1 Supported in part by UNAM-DGAPA Postdoctoral fellowship of the Institute of Mathematics, and by the DFG (grant CO 2359/1-1, Priority Programme SPP2026 “Geometry at Infinity”), and by a UKRI Future Leaders Fellowship [grant number MR/W01176X/1; PI: J Harvey].
2 Supported in part by the DFG (grant GA 2050 2-1, Priority Programme SPP2026 “Geometry at Infinity”).
1. Main results
@@
Singular Riemannian foliations generalize both isometric compact Lie group actions and Riemannian submersions, which induce decompositions into embedded submanifolds of lower dimension, and represent a generalized notion of symmetry on Riemannian manifolds [3, 10, 12, 19, 36]. Not all singular Riemannian foliations stem from Lie group actions (see, for example, [38]). Nevertheless, certain isometries of a Riemannian manifold $M$ with a singular Riemannian foliation $\mathcal{F}$ may induce residual symmetry by preserving the foliation’s leaves. These foliated isometries generate isometries of the leaf space $M/\mathcal{F}$. Our first main result is an analogue of the Myers–Steenrod theorem for the group of foliated isometries of a Riemannian manifold with a singular Riemannian foliation.
@@
    (a) $\dim(\operatorname{Isom}(M,\mathcal{F}))\leq\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2}.$
@@
    If equality holds in inequality (a), then $M$ is foliated-diffeomorphic to a fiber bundle $F\rightarrow M\to B$, where $B$ is diffeomorphic to $\mathbb{R}^{k}$, $\mathbb{R}P^{k}$, or $S^{k}$, and the fibers are round spheres or real projective spaces. Moreover, when equipped with the leaf-projection metric induced by $M$, the base space $B$ is isometric to $k$-dimensional Euclidean space, hyperbolic space with constant negative sectional curvature, a round real projective space, or a round sphere.
The leaf space $M/\mathcal{F}$ of a singular Riemannian foliation $(M,\mathcal{F})$ with closed leaves has a natural distance function that makes $M/\mathcal{F}$ a locally compact length space. The curvature of $M/\mathcal{F}$ is locally bounded from below in the triangle comparison sense, which links the geometry of the leaf space with that of the manifold. This fact plays a central role in proving Theorem A. For an isometric action of a compact Lie group $G$ on a complete Riemannian manifold $M$ with sectional curvature uniformly bounded from below by $k\in\mathbb{R}$, the orbit space $M/G$ equipped with the orbital distance function is an Alexandrov space with curvature bounded below by $k$. Imposing further conditions, such as positive or non-negative sectional curvature on $M$, leads to significant constraints on the manifold’s topology and has been an active research topic in Riemannian geometry [21, 22, 40]. Note that the orbit projection map $\pi\colon M\to M/G$ is a proper submetry, i.e., for every $p\in M$, any closed metric ball $B(p,r)$ of radius $r$ centered at $p$ maps onto the metric ball $B(\pi(p),r)$ in $M/G$. Submetries, introduced by Berestovskii in [5] as a metric generalization of Riemannian submersions, have been the focus of systematic study in metric geometry [4, 25, 27, 29].
Motivated by the preceding considerations, we also investigate the foliated isometries of the metric foliations $\mathcal{F}$ whose leaves are the fibers of proper submetries $\pi\colon X\to Y$ between Alexandrov spaces. By the work of Lytchak [29], the leaves of such foliations must be closed. We consider the group $\operatorname{Isom}(X,\mathcal{F})$ of foliated isometries and obtain the following analogue of Theorem A.
@@
Let $\pi\colon X\to Y$ be a proper submetry between connected Alexandrov spaces and set =$\mathcal{F}=\{\pi^{-1}(y)\mid y\in Y\}$. Then the following assertions hold:
@@
    If $X$ has dimension $n\geq 1$ and $Y$ has dimension $0\leqslant k\leqslant n$, then
    (b) $\dim(\operatorname{Isom}(X,\mathcal{F}))\leq\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2}.$
@@
    If equality holds in inequality (b), then $X$ is foliated-homeomorphic to a fiber bundle $F\rightarrow X\to Y$, where $Y$ is homeomorphic to $\mathbb{R}^{k}$, $\mathbb{R}P^{k}$, or $S^{k}$, and the fibers are isometric to a finite disjoint union og round spheres or projective spaces. Moreover, when equipped with the leaf-projection metric induced by $X$, the base space $Y$ is isometric to $k$-dimensional Euclidean space, hyperbolic space with constant negative sectional curvature, a round real projective space, or a round sphere.
@@
The upper bound for the dimension of the group of foliated isometries in Theorems A and B consists of two summands. The first summand, $k(k+1)/2$, bounds the dimension of the group of foliated isometries $f$ that do not leave all leaves invariant, i.e., there exists some leaf $L\in\mathcal{F}$ with $f(L)\not\subset L$. These are the isometries that descend to non-trivial isometries of the leaf space. The second summand, $(n-k)(n-k+1)/2$, bounds the dimension of the group of foliated isometries $h$ that leave the leaves invariant, i.e., $h(L)\subset L$ for any leaf $L\in\mathcal{F}$.
@@
Applying Theorem A to the trivial foliations =$\mathcal{F}=\{M\}$ or =$\mathcal{F}=\{{p}\mid p\in M\}$ consisting, respectively, of a single leaf or leaves that contain only one point, yields the classical upper bound $n(n+1)/2$ in the Riemannian Myers–Steenrod theorem. For Alexandrov spaces, Theorem B applied to the trivial submetries $\mathrm{id}\colon X\to X$ and $\pi\colon X\to\{\mathrm{pt}\}$ yields the bound $n(n+1)/2$ in the Myers–Steenrod theorem for Alexandrov spaces in [18].
@@
As in the Riemannian, Alexandrov, and $\mathrm{RCD}$ Myers–Steenrod theorems, the upper bound in Theorems A and B is sharp. Consider the Riemannian product $(S^{n-k}\times S^{k},g_{1}\times g_{2})$ of two round spheres $(S^{k},g_{1})$, $(S^{n-k},g_{2})$ of different radii, with $n>2k$, and the Riemannian foliation $\mathcal{F}$ whose leaves are $S^{n-k}$. Then $\operatorname{Isom}(M,g)=\operatorname{Isom}(S^{k},g_{1})\times\operatorname{Isom}(S^{n-k},g_{2})$, and thus $\operatorname{Isom}(M,\mathcal{F})=\operatorname{Isom}(M)$. This realizes the upper bound in Theorem A (a) and Theorem B (b).
@@
When equality holds in inequality (a), $M$ is a fiber bundle $F\to M\to B$ and we may assume that the leaves $F$ carry a metric of constant sectional curvature equal to one. The main theorems in [11] and [15] imply that, when $B$ is not diffeomorphic to $\mathbb{R}^{k}$, the bundle has a linear structure group. When $B$ is diffeomorphic to $\mathbb{R}^{k}$, the manifold $M$ is diffeomorphic to $\mathbb{R}P^{n-k}\times\mathbb{R}^{k}$ or $S^{n-k}\times\mathbb{R}^{k}$, since $\mathbb{R}^{k}$ is contractible. This last conclusion also holds for the case of Theorem B (iii) in the case when $Y$ is homeomorphic to $\mathbb{R}^{k}$.
@@
The map $\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M/\mathcal{F})$ is generally not surjective. However, if $M$ is an $n$-dimensional Euclidean vector space with a foliation $\mathcal{F}$ induced by a linear isometric action by a compact Lie group $G$, any isometry in $\operatorname{Isom}(M/G)_{o}$ lifts to a foliated isometry in $\operatorname{Isom}(M,G)$ [33]. Here, $\operatorname{Isom}(M/G)_{o}$ denotes the connected component containing the identity. The proof of Theorem B yields an upper bound on the dimension of the orbits of $G$ in terms of the dimension of $M/G$. Specifically, if $M/G$ has dimension $0<k<n$, then an orbit has dimension at most $(n-k)(n-k+1)/2$.
@@
We point out that in the proof of Theorem A, we only use the transnormal properties of the closed foliation. Nonetheless, we state our results for singular Riemannian foliations, as there are no known examples of transnormal systems that are not smooth foliations, and it is conjectured that any transnormal system is a singular Riemannian foliation (see [40, Final Remarks]).
@@
In the proof of Theorem A (i), we heavily rely on the assumption that the leaf space is Hausdorff. To prove item Theorem A (ii), we require the leaf space to be an Alexandrov space. Both conditions are satisfied if the singular Riemannian foliation has closed leaves. However, this requirement can be relaxed by asking that the leaves of the foliation be globally equidistant instead of just locally equidistant.
@@
We do not know if the upper bound on the dimesion stated in Theorem A (ii) holds for singular Riemannian foliations without closed leaves. Even if such a bound exists, we cannot obtain a rigidity conclusion as in Theorem A (iii). This is illustrated by the irrational flow on the $2$-dimensional flat torus: the group of foliated isometries has dimension two, which is the dimension of the isometry group of the $2$-dimensional flat torus. In contrast, when we assume that the leaves are closed and the singular Riemannian foliation is not trivial, Theorem A (iii) implies that the foliation is given by a circle bundle over the circle.
@@
In þB (iii), there are examples of foliations with disconnected fibers. Namely, consider the product of a round $2$-sphere with a round $3$-sphere, and conisder the $3$-sphere as leaves. As remarked above, the group of foliated isometries has maximal dimension. Now consider the antipodal action of $\mathbb{Z}_{2}$ on $S^{2}$. With this, we get a fibration of $S^{2}\times S^{3}$ over $\mathbb{R}P^{2}$ whose fibers are two disjoint copies of a round $S^{3}$. The difference with þA is that we ask the leaves of a singular Riemannian foliation to be connected.
@@
Acknowledgements.
We thank Alexander Lytchak and Marco Radeschi for helpful comments on a preliminary version of this article. We thank the organizers of the IV joint meeting of RSME and SMM, and the Universidad Politécnica de Valencia, for their hospitality while this manuscript was finished.
2. Preliminaries
In this section, we collect basic definitions and results on singular Riemannian foliations, Alexandrov spaces, and submetries we will use in the proof of our main theorems. We refer the reader to [2, 20, 8, 29, 27] for further basic results on theses subjects. We will assume all spaces to be connected, unless stated otherwise.
2.1. Singular Riemannian foliations
A singular Riemannian foliation $(M,\mathcal{F})$ on a complete Riemannian manifold $M$ is a partition of the manifold into a collection =$\mathcal{F}=\{L_{x}\mid x\in M\}$ of disjoint connected, complete, immersed submanifolds $L_{x}$, called leaves, satisfying the following conditions:

### v3

@@
[…]
Myers–Steenrod Theorems for Metric and Singular Riemannian Foliations
[…]
Myers–Steenrod Theorems for Metric
and Singular Riemannian Foliations
[…]
Diego CORRO ab and Fernando GALAZ-GARCÍA c
[…]
D. Corro and F. Galaz-García
[…]
a) Fakultät für Mathematik, Karlsruher Institut für Technologie, Germany
[…]
Received November 04, 2024, in final form December 01, 2025; Published online December 16, 2025
[…]
We prove that the group of isometries preserving a metric foliation on a closed Alexandrov space $X$ is a closed subgroup of the isometry group of $X$. We obtain a sharp upper bound for the dimension of this subgroup and show that, when equality holds, the foliations that realize this upper bound are induced by fiber bundles whose fibers are round spheres or projective spaces. As a corollary, singular Riemannian foliations that realize the upper bound are induced by smooth fiber bundles whose fibers are round spheres or projective spaces.
[…]
1 Main results
@@
Singular Riemannian foliations generalize both isometric compact Lie group actions and Riemannian submersions, which induce decompositions into embedded submanifolds of lower dimension, and represent a generalized notion of symmetry on Riemannian manifolds [3, 13, 16, 24, 46]. Not all singular Riemannian foliations stem from Lie group actions (see, for example, [48]). Nevertheless, certain isometries of a Riemannian manifold $M$ with a singular Riemannian foliation $\mathcal{F}$ may induce residual symmetry by preserving the foliation’s leaves. These foliated isometries generate isometries of the leaf space $M/\mathcal{F}$. Our first main result is an analog of the Myers–Steenrod theorem for the group of foliated isometries of a Riemannian manifold with a singular Riemannian foliation.
@@ 2. $(ii)$
    $\dim(\operatorname{Isom}(M,\mathcal{F}))\leq\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2}.$ (a)
@@ 3. $(iii)$
    If equality holds in inequality (a), then $M$ is foliated-diffeomorphic to a fiber bundle $F\to M\to B$, where $B$ is diffeomorphic to $\mathbb{R}^{k}$, $\mathbb{R}P^{k}$, or $S^{k}$, and the fibers are round spheres or real projective spaces, Euclidean space, or a hyperbolic space. Moreover, when equipped with the leaf-projection metric induced by $M$, the base space $B$ is isometric to $k$-dimensional Euclidean space, hyperbolic space with constant negative sectional curvature, a round real projective space, or a round sphere.
The leaf space $M/\mathcal{F}$ of a singular Riemannian foliation $(M,\mathcal{F})$ with closed leaves has a natural distance function that makes $M/\mathcal{F}$ a locally compact length space. The curvature of $M/\mathcal{F}$ is locally bounded from below in the triangle comparison sense, which links the geometry of the leaf space with that of the manifold. This fact plays a central role in proving Theorem A. For an isometric action of a compact Lie group $G$ on a complete Riemannian manifold $M$ with sectional curvature uniformly bounded from below by $k\in\mathbb{R}$, the orbit space $M/G$ equipped with the orbital distance function is an Alexandrov space with curvature bounded below by $k$. Imposing further conditions, such as positive or non-negative sectional curvature on $M$, leads to significant constraints on the manifold’s topology and has been an active research topic in Riemannian geometry [26, 27, 51]. Note that the orbit projection map $\pi\colon M\to M/G$ is a proper submetry, i.e., for every $p\in M$, any closed metric ball $B(p,r)$ of radius $r$ centered at $p$ maps onto the metric ball $B(\pi(p),r)$ in $M/G$. Submetries, introduced by Berestovskii in [5] as a metric generalization of Riemannian submersions, have been the focus of systematic study in metric geometry [8, 30, 32, 36, 37, 39].
Motivated by the preceding considerations, we also investigate the foliated isometries of the metric foliations $\mathcal{F}$ whose leaves are the fibers of a submetry $\pi\colon X\to Y$ between Alexandrov spaces. Each leaf $\pi^{-1}(y)$ is closed, since it is the inverse image of the closed set $\{y\}$ under the continuous map $\pi$. We consider the group $\operatorname{Isom}(X,\mathcal{F})$ of foliated isometries and obtain the following analog of Theorem A. For parts (ii)–(iii), we assume $\pi$ is proper, in which case every fiber is compact.
@@
Let $\pi\colon X\to Y$ be a submetry between connected Alexandrov spaces and set $\mathcal{F}=\big\{\pi^{-1}(y)\mid y\in Y\big\}$. Then the following assertions hold:
@@ 1. $(ii)$
    If $X$ has dimension $n\geq 1$ and $Y$ has dimension $0\leq k\leq n$, then
    $\dim(\operatorname{Isom}(X,\mathcal{F}))\leq\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2}.$ (b)
@@ 2. $(iii)$
    If equality holds in inequality (b), then $\pi$ is the composition of a submetry $\tilde{\pi}\colon X\to Z$ with connected fibers, and a submetry $\pi_{D}\colon Z\to Y$ whose fibers are finite discrete spaces. The spaces $X$, $Y$, $Z$ are Riemannian manifolds, $X$ is homogeneous, and the submetry $\tilde{\pi}\colon{X\to Z}$ is a smooth Riemannian submersion. Moreover, the conclusions of Theorem A(iii)(iii) hold for the foliation induced by $\tilde{\pi}$. In general the fibers of $\pi$ are a finite disjoint union of round spheres or projective spaces, the base space $Y$ is isometric to the $k$-dimensional Euclidean space, hyperbolic space with constant negative sectional curvature, a round real projective space, or a round sphere. More specifically, when $Y$ is homeomorphic to $\mathbb{R}^{k}$ for $k\geq 0$, or $Y$ is homeomorphic to $S^{k}$ for $k\geq 2$, then the fibers of $\pi$ are connected. When $Y$ is homeomorphic to $\mathbb{R}P^{k}$ for $k\geq 2$, the fibers of $\pi$ have at most two connected components. When $Y$ is homeomorphic to $S^{1}$, then the fibers of $\pi$ can have $m\geq 1$ connected components.
@@
The upper bound for the dimension of the group of foliated isometries in Theorems A and B consists of two summands. The first summand, $k(k+1)/2$, bounds the dimension of the image of the Lie group morphism $\Psi\colon\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M/\mathcal{F})$ induced by the projection $\pi$ (see Section 2.7). This image consists of the isometries of $M$ that descend to non-trivial isometries of the leaf space. The second summand, (nk)(nk+1)/2(n-k)(n-k+1)/2, bounds the dimension of the group of foliated isometries $h$ that leave the leaves invariant, i.e., $h(L)\subset L$ for any leaf $L\in\mathcal{F}$.
@@
Applying Theorem A to the trivial foliations $\mathcal{F}=\{M\}$ or $\mathcal{F}=\{{p}\mid p\in M\}$ consisting, respectively, of a single leaf or leaves that contain only one point, yields the classical upper bound n(n+1)/2{n(n+1)/2} in the Riemannian Myers–Steenrod theorem. For Alexandrov spaces, Theorem B applied to the trivial submetries $\mathrm{id}\colon X\to X$ and $\pi\colon X\to\{\mathrm{pt}\}$ yields the bound $n(n+1)/2$ in the Myers–Steenrod theorem for Alexandrov spaces in [23].
@@
As in the Myers–Steenrod theorem for the setting of Riemannian, Alexandrov, and $\mathrm{RCD}$-spaces, the upper bound in Theorems A and B is sharp. Let $(M,g)$ be the Riemannian product $(N\times P,g_{1}\times g_{2})$} of two Riemannian manifolds $\bigl(N^{n-k},g_{1}\bigr)$, $\bigl(P^{k},g_{2}\bigr)$ isometric to round spheres or round projective spaces with $k,n-k\geq 1$, and the Riemannian foliation $\mathcal{F}$ whose leaves are $N$. Then $\dim(\operatorname{Isom}(M,g))=\dim(\operatorname{Isom}(N,g_{1}))+\dim(\operatorname{Isom}(P,g_{2}))$ (see [19, Corollary 1]), and thus $\dim(\operatorname{Isom}(M,\mathcal{F}))=\dim(\operatorname{Isom}(M))$. This realizes the upper bound in Theorem A (a) and Theorem B (b).
As a nontrivial bundle example, consider the Hopf fibration $\mathcal{F}$ given by the orbits of the scalar multiplication action of $S^{1}\subset\mathbb{C}$ on the unit sphere $S^{3}\subset\mathbb{C}^{2}$. Since this $S^{1}$-action commutes with the standard $\mathrm{U}(2)$-action on $S^{3}$, we have $\mathrm{U}(2)\subset\operatorname{Isom}\bigl(S^{3},\mathcal{F}\bigr)$. By Theorem A(ii)(ii),
$\dim\mathrm{U}(2)=4\leq\dim\bigl(\operatorname{Isom}\bigl(S^{3},\mathcal{F}\bigr)\bigr)\leq 4.$
Hence, $\operatorname{Isom}\bigl(S^{3},\mathcal{F}\bigr)$ has maximal dimension while the foliation is given by a nontrivial fiber bundle.
@@
When equality holds in inequality (a), $M$ is a fiber bundle $F\to M\to B$ and we may assume that the leaves $F$ carry a metric of constant sectional curvature equal to one. The main theorems in [14, 20] imply that, when $B$ is not diffeomorphic to $\mathbb{R}^{k}$, the bundle has a linear structure group. In particular, when the leaves are diffeomorphic to $S^{n-k}$, the classification of such bundles is the same as the classification of vector bundles over $B$. In the case when the leaves are diffeomorphic to $\mathbb{R}P^{n-k}$, in some cases depending on the values of $k$, we can show that the conditions of [4, Theorem 1] hold, and thus the classification reduces to the classification of vector bundles over $B$. In general when the leaves are diffeomorphic to $\mathbb{R}P^{n-k}$, by [14], the structure group is the so-called projective orthogonal group $\mathrm{PO}(n-k)=\mathrm{O}(n-k)/\{\mathrm{Id},-\mathrm{Id}\}$, and thus the classification corresponds to $B,B\mathrm{PO}(n-k)]$, the collection of maps from $B$ to the classifying space $B\mathrm{PO}(n-k)$ up to homotopy.
When $B$ is diffeomorphic to $\mathbb{R}^{k}$, the manifold $M$ is diffeomorphic to $\mathbb{R}P^{n-k}\times\mathbb{R}^{k}$ or $S^{n-k}\times\mathbb{R}^{k}$, since $\mathbb{R}^{k}$ is contractible. Observe that in Theorem B(iii)(iii) in the case when $Y$ is homeomorphic to $\mathbb{R}^{k}$, we conclude that $X$ is homeomorphic to $S^{n-k}\times\mathbb{R}^{k}$ or $\mathbb{R}P^{n-k}\times\mathbb{R}^{k}$.
@@
The map $\Psi\colon\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M/\mathcal{F})$ is generally not surjective. However, if $M$ is an $n$-dimensional Euclidean vector space with a foliation $\mathcal{F}$ induced by a linear isometric action by a compact Lie group $G$, any isometry in $\operatorname{Isom}(M/G)^{0}$ lifts to a foliated isometry in $\operatorname{Isom}(M,G)$ [43]. Here, $\operatorname{Isom}(M/G)^{0}$ denotes the connected component containing the identity. The proof of Theorem B yields an upper bound on the dimension of the orbits of $G$ in terms of the dimension of $M/G$. Specifically, if $M/G$ has dimension $0<k<n$, then an orbit has dimension at most (nk)(nk+1)/2{(n-k)(n-k+1)/2}.
@@
We point out that in the proof of Theorem A, we only use the transnormal properties of the closed foliation. Nonetheless, we state our results for singular Riemannian foliations, as there are no known examples of transnormal systems that are not smooth foliations, and it is conjectured that any transnormal system is a singular Riemannian foliation (see [51, Final remarks]). See also the work of Lytchak and Wilking [39] for the smoothness of submetries between Riemannian manifolds.
@@
In the proof of Theorems A(i)(i) and B(i)(i), we crucially rely on the assumption that the leaf and fiber spaces are Hausdorff (see, for example, the proof of Proposition 2.23). To prove Theorem A(ii)(ii), we require the leaf space to be an Alexandrov space. Both conditions are satisfied if the singular Riemannian foliation has closed leaves. However, this requirement can be relaxed by asking that the leaves of the foliation be globally equidistant instead of just locally equidistant.
@@
We do not know if the upper bound on the dimension stated in Theorem A(ii)(ii) holds for singular Riemannian foliations without closed leaves. Even if such a bound exists, we cannot obtain a rigidity conclusion as in Theorem A(iii)(iii). This is illustrated by the irrational flow on the $2$-dimensional flat torus: the group of foliated isometries has dimension two, which is the dimension of the isometry group of the $2$-dimensional flat torus. In contrast, when we assume that the leaves are closed and the singular Riemannian foliation is not trivial, Theorem A(iii)(iii) implies that the foliation is given by a circle bundle over the circle.
@@
In Theorem B(iii)(iii), there are examples of foliations with disconnected fibers. Namely, consider the product of a round $2$-sphere with a round $3$-sphere, and consider the $3$-sphere as leaves. As remarked above, the group of foliated isometries has maximal dimension. Now consider the antipodal action of $\mathbb{Z}_{2}$ on $S^{2}$. With this, we get a fibration of $S^{2}\times S^{3}$ over $\mathbb{R}P^{2}$ whose fibers are two disjoint copies of a round $S^{3}$. The difference with Theorem A is that we ask the leaves of a singular Riemannian foliation to be connected.
Remark 1.5.
We note a natural classification problem: classify, up to foliated isometry, all pairs $(X,\mathcal{F})$ with $\operatorname{Isom}(X,\mathcal{F})$ of maximal dimension. Our results provide initial structure constraints for this question.
@@
2 Preliminaries
In this section, we collect basic definitions and results on singular Riemannian foliations, Alexandrov spaces, and submetries we will use in the proof of our main theorems. We refer the reader to [2, 11, 25, 32, 36] for further basic results on these subjects. We will assume all spaces are connected, unless stated otherwise.
2.1 Singular Riemannian foliations
A singular Riemannian foliation $(M,\mathcal{F})$ on a complete Riemannian manifold $M$ is a partition of the manifold into a collection $\mathcal{F}=\{L_{x}\mid x\in M\}$ of disjoint connected, complete, immersed submanifolds $L_{x}$, called leaves, satisfying the following conditions: